What is the value of wave funct at the surface of an infinite potential is
1)0
2) infinite
3)1
1
Expert's answer
2019-04-11T09:51:47-0400

The wave function has the form


"\varPsi_n=B\sin{\frac {n\pi x} {L}} (1)"

For infinite potential we have:


"L\rightarrow \infty (2)"

We put (2) in(1) and got:


"\varPsi_n=0"

Answer:

1)0